OpenStudy (squirrels):

What is a possible value for the missing term of the geometric sequence? 37, ___, 925...? a. -185 b. -4,625 c. -925 d. 481 The answer is obviously d. correct? I feel like this is a trick question, not to sure.

OpenStudy (anonymous):

why obviously???

OpenStudy (squirrels):

I figured that it couldn't be a negative lol

OpenStudy (anonymous):

Its A

OpenStudy (anonymous):

Nah it cld be

OpenStudy (squirrels):

tell me how c:<

OpenStudy (anonymous):

I just taught this today

OpenStudy (anonymous):

Ok we know that we start off with our first term, called "a" In this case a=37

OpenStudy (squirrels):

Indeed

OpenStudy (anonymous):

Then subsequently we keep multiplying the previous term by a constant We are not given the constant

OpenStudy (anonymous):

But the constant is known as r

OpenStudy (anonymous):

so our first term is \[ \{ a, a*r, a*r*r, a*r*r*r, ...\}\] which in other words is \[ \{a,ar,ar^2,ar^3,ar^4,,,\}\]

OpenStudy (anonymous):

You familiar with this?

OpenStudy (squirrels):

Hmm I'm just learning this, but I get what you're saying.

OpenStudy (anonymous):

SO in our case we have the first term is 37 second is unknown and third term is 925 so a=37, ar=37*r and ar^2=925

OpenStudy (squirrels):

Ok I get what you're saying..

OpenStudy (anonymous):

so ar^2=925 we know that a=37 so \(37*r^2=925\) \[ r^2=\frac{925}{37}\]

OpenStudy (anonymous):

so \[r^2=25\] \[\sqrt{r^2}=\sqrt{25}\] r=5

OpenStudy (anonymous):

So now we found out that r=5

OpenStudy (anonymous):

So its easy to find the terms

OpenStudy (anonymous):

First term is "a" which equals 37 Second term is a*r=37*5=q85 and the third term is ar^2=37(5)^2=925

OpenStudy (anonymous):

So we see that our second term is 185

OpenStudy (anonymous):

Basically really r is really -5 a=37 a*r=37(-5)=-185 a*r*r=37(-5)(-5)=925

OpenStudy (anonymous):

Is this an online course?

OpenStudy (squirrels):

Yes mam, it is, and thank you Rach c: but one more question. Why is it negative?

OpenStudy (anonymous):

GONNAAA KILLL YOU :P

OpenStudy (squirrels):

Lmao cx is there a reason though? hahaha

OpenStudy (anonymous):

Ok i was showing you that the answer can be negative or positive in reality

OpenStudy (squirrels):

ohhhh i see now lol thanks hahah cx

OpenStudy (anonymous):

Ya rightttt :PPP I can show you an example if you would like

OpenStudy (squirrels):

lol no I'm serious hahaha, i understood your method! I just didn't know why it was negative lol

OpenStudy (anonymous):

Okkk lemme show u an example real quick

OpenStudy (squirrels):

Of course :3

OpenStudy (squirrels):

And thanks c:

OpenStudy (anonymous):

Ok so lets take this example where our starting term a=37 and the multiplying constant is -5 Please list the first 5 terms of the sequence

OpenStudy (squirrels):

Ok so 37, 185, 925, 4625, 23125 correct?

OpenStudy (anonymous):

Nahhh remember its -5

OpenStudy (anonymous):

Take the sign into account

OpenStudy (squirrels):

My bad, 37, -185, 925, -4625, 23125, is this better?

OpenStudy (anonymous):

Yupppp Now refer back to the question above

OpenStudy (squirrels):

So the answer is a. :3

OpenStudy (squirrels):

Is that what you're asking me?

OpenStudy (anonymous):

Yaaaaaa

OpenStudy (anonymous):

The answer can technically be 185 or -185 but since they only gave the option of -185 we chose that

OpenStudy (squirrels):

Ok i see whatchu mean B) Thanks a lot Rachel c:

OpenStudy (anonymous):

Okkk Im out for now but i may be back later

OpenStudy (squirrels):

Ok Love You

OpenStudy (anonymous):

Awwwww ♥

OpenStudy (squirrels):

Lol cx